Shift summary for eve. 11/17/99

Ed Blucher (blucher@hep.uchicago.edu)
Thu, 18 Nov 1999 00:07:38 -0600


Shift Summary for Eve Shift, 11/17/99

On Shift: Rick Ford, Ed Blucher

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...):0.5

Runs ended this shift:

14995 Laser scan

========================================================================
Overview of shift:
Shutdown recovery

========================================================================
Description of Shift:

16:50 Standard spill signal restored.

17:00 JW,EB,JH Swap CsI divider 2893 to fix lousy dynode signal. Signal
still lousy after swap. Channels 2893 and 2871 show large negative
pmt gain drift (-25%) after swap. It's possible that fibers were
bumped, so we'll wait to see response with electrons before adjusting
gains.

CsI LV sector 9 has a flakey sequencer; it should be checked during
the next long access.

CsI channel 51 reads 75V when demand is 1140V; it appears that voltage
sensing is screwed up for this channel.

18:30 PSH changed laser delay from 8ns to 14ns.

20:20 CsI Sector 8 LV alarm on "-5.2V" supply. Readback is ~5.3 V compared to
nominal 6V. After consulting with BobT, we make an access to check
voltage. We are unable adjust supply to go above 5.7V.
BT and EB replace -5.2V supply.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Complete muon run.

Check E/p for CsI channels 2893 and 2871.